OpenStudy (star):

the limit as x-> 0 for cos(2x)-1 / cos(3x)-1

myininaya (myininaya):

ok if you plug in 0 you get 0/0 so you can use l'hospital's rule -2sin(2x)/-3sin(3x) give me just a sec i'm going to use the equation thingy

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{2}{3} \frac{\sin(2x)}{2x}\frac{\sin(3x)}{3x}6x^2\]

myininaya (myininaya):

2/3(1)(1)(6)(0)=0

OpenStudy (star):

it says the answer is 4/9

myininaya (myininaya):

\[\lim_{x \rightarrow 0} \frac{\cos(2x)-1}{\cos(3x)-1}\]

myininaya (myininaya):

is that right?

OpenStudy (star):

yep

myininaya (myininaya):

ok let me see

myininaya (myininaya):

oops i see a mistake

myininaya (myininaya):

i put sin(3x) in the numerator do you see that?

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{2}{3}\frac{\sin2x}{2x}\frac{3x}{\sin(3x)}\frac{2x}{3x}\]

myininaya (myininaya):

2/3*1*1*2/3=4/9

myininaya (myininaya):

got it? do you remember what sinx/x goes to as x goes to 0?